the coefficient of x49 in the expansion (x-1)()(x-1/2)(x-1/22)........(x-1/249) is???

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given expression is
=(x-1)(x-1/2)(x-1/22)........(x-1/249)
=x50-((1/2)0+(1/2)1+.......+(1/2)49)x49+.......
so coefficient of x49=(1/2)0+(1/2)1+.......+(1/2)49
As there are in GP with a common ratio r=(1/2) with a=1
coefficient of x49=1(((1/2)50-1)/(1/2)-1)
=(250-1)/250